Week of 181105

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Kate (chair, WLCG, DB), Maarten (ALICE, WLCG), Vladimir (LHCb), Ivan (ATLAS), Olga (computing), Paul (storage), Vincent (security), Marian (networks)
  • remote: John (RAL), Xavier (KIT), David (IN2P3), Di (TRIUMF), Christoph (CMS), Christian (NDGF), Xin (BNL), Dave M (FNAL), Eygene R (NRC-KI)

Experiments round table:

  • CMS reports ( raw view) -
    • Disk space situation improved
      • Some cleaning performed
      • Got some disk space contribution increase from sites thanks to early availability of 2019 pledges - Thanks!
    • Good CPU utilization again since ~end of last week, ~160k cores for production and ~50k cores for analysis
    • Issues with RAL FTS3 service GGUS:138002 (likely IPv6 related)

  • ALICE -
    • NTR

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF: nc
  • EGI: nc
  • FNAL: NTR
  • IN2P3:
    • Bad manipulation on Sunday results in decreasing number of jobs for LHCb. Apologies for this. Error is now fixed and situation is coming back to normal.
    • Still transfer problem are under investigation on network side. This has an impact for ATLAS also.
  • JINR: nc
  • KISTI: nc
  • KIT:
    • The downtime for atlassrm-fzk.gridka.de, that was planned for this Wednesday, has been canceled this morning. We wanted to first understand and resolve the IPv6 connectivity issues, before potentially exposing the ATLAS storage element to the same risks. We want to propose CW 47 (Nov 19th-23rd) as next possible date for a downtime to ATLAS.
    • Around noon today, the IPv6 connectivity issues between GridKa and RAL / Helsinki were resolved (routing problem regarding LHCOPN at CERN).
    • A couple of dCache service components crashed with out-of-memory errors on Friday and Sunday, though that seemingly had no measurable impact on the CMS availability. Our on-call engineer restarted the Java virtual machines swiftly. We'll have to contact dCache.org to ask for assistance in analysing the heap storage dump, to find out what had consumed the memory.
  • NDGF: Triolith decommissioned. Downtime ended. Bluegrass soon ready to replace it.
  • NL-T1: nc
  • NRC-KI:
    • Library upgrade was smooth, it's back in production since Wednesday, 30.10.2018.
    • There is an ongoing process of requesting GEANT (together with GARR) to expand our current peering link in Amsterdam from 10G to 20G as the current link is saturated and GARR wants the R&E traffic (as opposed to the LHCONE) from Tier-2 centers in Italy to be passed via GEANT and not our current direct peering via CERNlight (GARR has 100G to GEANT and only 10G to CERNlight).
  • OSG: nc
  • PIC: nc
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services: Post-meeting note: (OTG:0046717) Intervention planned for next week, expected VOMS service downtime.
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring: nc
  • MW Officer: nc
  • Networks: NTR
  • Security: EGI found an issue with their security monitoring, which has been fixed.

AOB:

Edit | Attach | Watch | Print version | History: r17 < r16 < r15 < r14 < r13 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r17 - 2018-11-06 - OlgaVladimirovnaDatskova
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback